Hydrogen atom in states of high quantum number have been created in the laboratory and observed in space.
(a) Find the quantum number of the bohr orbit in a H-atom whose radius is 0.0100mm?
(b) What is the energy of H-atom in this state?

Grade:11

2 Answers

Vikas TU
7 years ago
a) Bohr radii is given by:
 
r = 0.529*(n^2/z)
0.0100 *10^7 A= 0.529*n^2 {For H, z= 1, A is angstorm}
n = root(189035.91)
= 434.7
which can be taken as 435
.i.e n = 435
 
b) For energy
E = -13.6*n^2 = -13.6*435*435 = -2573460 J
OR
2459.925 KJ 
 
 
Neha
3 years ago
The quantum number given in this answer n=435 is correct but there is something wrong in energy .So, energy is En=-13.6z^2/n^2 
  -13.6×1^2/435*435 = -0.7187×10^-4 ev